Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary.

a = ?, b = 5, c = 8
A. 6.0
B. 6.1
C. 6.2
D. 6.3
Answer #1
a^2 + b^2 = c^2
answered by: Molly :)
Answer #2
a^2 + b^2 = c^2

a^2 + 5^2 = 8^2

a^2 + 25 = 64

a^2 = 64 - 25

a^2 = 39

a = 6.2449 = 6.2
answered by: MSM